1.)A 20lbm object is located 300 ft above the ground. What is the potential energy of the object in lbf-ft, BTU, and cal?


2.)A mole of a certain ideal gas was inside a container. The gas has a temperature of 70°F and a pressure of 20 psi. Find the volume of the gas in m^3.


1
Expert's answer
2021-12-16T13:13:02-0500

Momentum after release of steam P.F.f 

Pf = mv + m'v'       (1)

Where mass of empty launch steam m=7500 kg

m' is mass of satellite , m'=250 kg

v is final velocity of launch steam after releasing satellite, v=900 m/s

v' is final velocity of the satellite after release

Putting all the values in equation (1), we get

Pf =(7500 kg)(900 m/s) + (250 kg)v'

Pf = 6.750×10kg•m/s + (250 kg)v'


Now applying conservation of momentum

Momentum before release of satellite = momentum after release of satellite

Pi = Pf

7.750×106 kg•m/s = 6.750×10kg•m/s + (250 kg)v'

(7.750×106 - 6.750×10)kg•m/s= (250 kg)v'

106 m/s = 250v'

v' = (106/250) m/s

v' = 4000 m/s